Not sure if this is the right topic but here is the problem:
Find the equation of the line containing the point p(-1,3) and perpendicular to the graph of 2x+4y=-1.

Answer by addingup(3677)

We use the formula for the standard slope-intercept: y= mx+b, where m is the slope and b the y-intercept (the point at which the slope line touches the vertical y line. Y is always the vertical, x is always the horizontal):
Line perpendicular to the graph of:
2x + 4y = -1
Divide both sides by 2x+4:
y = -(1/2)x - 1/4 - Green line on graph
----------------------------------
The slope of the line is m= 2 (negative reciprocal of -1/2. If you don't remember how to do this: To find the negative reciprocal, flip the number and then if it's positive make it negative -- and if it's negative make it positive)
---------------------------
Now let's use the ordered pair (-1,3) to solve for b:
y = 2x+ b
3 = -2 + b
5 = b
y = 2x + 5 - Blue line on graph